# Quantum Langevin Dynamics for Optimization

### v1

@@ 2. $f(T)\geq T$ for $\forall T>0$,
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}+\eta\cdot\left(\braket{H}_{t}-2kT(t)\right)\leq 0.$ (75)
Defining a function $h(t)=(\braket{H}_{t}-2kf(T(t)))e^{\eta t}$, its time derivative can be calculated as
$\frac{\mathrm{d}h(t)}{\mathrm{d}t}=$ $\left(\frac{\mathrm{d}}{\mathrm{d}t}\braket{H}_{t}-2k\frac{\mathrm{d}f}{\mathrm{d}T}\frac{\mathrm{d}T}{\mathrm{d}t}\right)e^{\eta t}$
$+(\braket{H}_{t}-2kf(T(t)))\eta e^{\eta t}.$ (76)
@@
$\left(\frac{\mathrm{d}}{\mathrm{d}t}\braket{H}_{t}-2k\frac{\mathrm{d}f}{\mathrm{d}T}\frac{\mathrm{d}T}{\mathrm{d}t}\right)+\eta(\braket{H}_{t}-2kf(T(t)))$ (78)
$\leq$ $\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}+\eta\cdot\left(\braket{H}_{t}-2kT(t)\right)\leq 0,$ (79)
@@
Then $h(t)\leq h(0)$ implies $\braket{H}_{t}-2kf(T(t))\leq O(e^{-\eta t})$. Using Eq. (19), we get $\braket{V}_{t}\leq\braket{H}_{t}\leq 2kf(T(t))+O(e^{-\eta t})$.
@@
V.1.4 Three time-dependent parameters $\eta(t),\hbar(t),T(t)$
@@
whose proof, similar to the proof of Corollaries 1, 2 and 3, is omitted here.
V.2 Objective functions
@@
Ref. [32] also explains how to categorize different objective functions. The category “Ridges or Valleys” is featured by small deep sections that divide the domain. “Basin” is featured by flatness near the global minimum of function values. “Flat” is featured by its flatness at elevated function values. “Studded” is identified by a foundational shape that exhibits high-frequency perturbations. The category “Simple” represents functions that can be solved efficiently by standard gradient descent algorithms. The functions in each category are listed in Table 1, whose precise formulations can be found in Appendix A.2.
@@
V.3 Time-dependent Quantum Langevin Dynamics
In the previous experiments, we analyze roles of $T,\eta,\hbar$ in QLD, which also lead to the question why QLD would have a better performance if these parameters are time-dependent, as different ranges of parameters have different effect in different phases of the optimization process. At the beginning of the evolution, we set a high temperature to prompt the system to escape from local minimum. Then, we reduce the temperature to cool the system, allowing it to reach a more optimal, lower energy state. The discussion concerning $\hbar$ is also similar. At the beginning of the evolution, larger $\hbar$ can accelerate the rate of quantum tunneling, helping system escape from local minimum. After seeking the global minimum, smaller $\hbar$, weakening the quantum tunneling effect, can prevent system from tunneling back to local minimum. If $\hbar$ and $T$ decrease at the same time, $\hbar$ must decrease faster than $T$ as the effect of tunneling back to local minimum is very significant at a mild temperature, termed as “quantum anti-tunneling”. This could be due to the large value of $\hbar$, which violates the approximation $\hbar\Omega\ll kT$ as stated in Corollaries 1, 2 and 3. This violation could trigger the “quantum anti-tunneling” phenomenon. Therefore, $\hbar(t)$ needs to decrease faster than $T(t)$ as Eq. (89) and Eq. (90). This might ensure that the system can concentrate on the global minimum after escaping from local minimum.
@@ V.3 Time-dependent Quantum Langevin Dynamics
Same finite difference method and implicit Euler method as Section IV.2 are used to numerically solve the evolution of system. The only difference is that time-dependent $\hbar(t)$ and $T(t)$ make $\hat{\mathcal{L}}$ of Eq. (96) be a time-dependent matrix ^$\hat{\mathcal{L}}(t)$. Dirichlet boundary condition and approximately uniform distribution are still kept here.
@@ V.4 Quantum algorithms
V.4.1 Quantum Hamiltonian Descent (QHD)
@@
Then, we discretize the space into grids similar to Section IV.2 and use symplectic leapfrog scheme [50] to evolve this Hamiltonian system.
@@
V.4.2 Quantum Adiabatic Algorithm (QAA)
@@ V.5 Classical algorithms
V.5.1 Stochastic Gradient Descent (SGD)
@@
V.5.2 Nesterov’s accelerated gradient descent (NAGD)
@@ VI Discussion and Conclusions
However, there are several issues that necessitate further investigation. Theoretically, we believe that both the advantages and shortcomings of QLD can be interpreted more thoroughly. In particular, due to the complex formula of the quantum Langevin dynamics, we only provide the convergence of QLD under the approximations of $\hbar\Omega\ll kT$ and $T\to 0$. Since the limit $\hbar\to 0$ clearly gives a transition to the classical Langevin equation, where all commutators vanish in this limit, in a sense, we might just prove the convergence of QLD in the classical limit, and use its quantum effect in the process of convergence. Whether the convergence still exists without these approximations remains further discussion with more meticulous mathematical and physical analysis. Numerically, since the simulation in the one-dimensional landscape alone has consumed a considerable amount of our computational resources, numerical experiments in high dimension requires more efficient numerical methods and more powerful classical computers. This also demonstrates the necessity of analog (or digital) quantum computers.
Acknowledgements.
We thank Jiaqi Leng, Hao Wu, Shuohan Zhang, and Zhennan Zhou for helpful discussions. Z. C. and Y. L. acknowledge research funding from Tsien Excellence in Engineering Program of Tsinghua University and Tsinghua University Initiative Scientific Research Program. T. L. was supported by the National Natural Science Foundation of China (Grant Numbers 62372006 and 92365117), and a startup fund from Peking University.
@@ A.1 Implementation details of numerical experiments for QLD
Because objective function reaches positive infinity at the boundary of domain in most cases, it is reasonable to use Dirichlet boundary condition in our numerical experiments. However, other efforts are still needed to avoid the instability of numerical calculation. In order to maintain the stability at the boundary, we set initial probability distribution to be an approximately uniform distribution with form
@@ A.2 The precise formulations of objective functions
We choose seven objective functions with different features in Table 1 to test the performance of different algorithms, as discussed in Section V.2. The precise formulations and global minimums $x^{*}$ of these objective functions defined in $x\in[-1,1]$ are listed below,
@@ A.3 Analysis of numerical results
The results of Michalewicz function are shown in Section V.6. Due to space constraints, the numerical results of rest six objective functions are shown here. The success probabilities of six objective functions for different algorithms at different (effective) evolution times are shown in Fig. 20. The probability distributions of six objective functions for different algorithms at different (effective) evolution times are shown in Figs. 21, 22, 23, 24, 25 and 26.
Based on the observations of Fig. 20, QLD exhibits much higher success probability across most objective functions than other algorithms. Specifically, the phenomena of strong quantum tunneling and thermal effects equip QLD with superior capabilities to escape from local minima encircled by numerous shallow wells at the beginning of evolution. This advantage enables QLD to achieve nearly 100% success probability in the Michalewicz, Alpine 1, Bohachevsky 2, and Deflected Corrugated Spring (DCS) functions which own several not-so-deep wells. However, the persistence of quantum noise and thermal effects towards the end of the evolution process hinder QLD’s ability to focus on the global minimum. Consequently, QLD performs poorly on the “Basin” function (Csendes), characterized by a flat region near the global minimum. The observations of dynamical processes in these experiments further confirms the three-phase explanation of QLD, which is shown at the end of Section V.6.
@@ B.1 Quantum query complexity for QLD
We can divide Hamiltonian into two parts $H=E_{k}+V$. Then by finite difference method mentioned in Section IV.2, it is very easy to present $E_{k}$ and $V$ in the matrix forms. Consider discretizing the domain into $n$ grids along each axis, $E_{k}=-\frac{\hbar^{2}\mathbf{L}}{2m\Delta x^{2}}$, where $\mathbf{L}$ is the Laplacian matrix associated to the central difference method and $\Delta x$ is the width of discretizing grid [33], and $V=\text{diag}(V(x_{1}),\ldots,V(x_{n}))$ is a diagonal matrix by discretizing $V(x)$ in the grid, which is given by the oracle $O_{V}$ [32].
@@
Definition 1 (Block-encoding, Definition 43 of Ref. [77]).
@@
Definition 2 (State preparation pair, Definition 51 of Ref. [77]).
@@
Proposition 3 (Linear combination of block-encoded matrices, Proposition 5 of Ref. [78]).
@@
$\tilde{W}_{A}=$ $\big{(}P_{L}^{\dagger}\otimes I_{a+q}\big{)}\left(\sum_{j=0}^{m-1}\ket{j}\bra{j}\otimes(I_{a-a_{j}}\otimes U_{j})\right.$
$\left.+\sum_{j=m}^{2^{b}-1}\ket{j}\bra{j}\otimes I_{a+q}\right)\big{(}P_{R}\otimes I_{a+q}\big{)}$
@@
Proposition 4 (Complexity for Lindblad simulation, Theorem 11 of Ref. [79]).
@@ B.2.3 Classical query complexity for SGD
For SGD, we assume that the classical evaluation oracle $O_{V}^{cl}$ allows us to query the function value of $V$ in a local way, so the total number of queries to $O_{V}^{cl}$ is at least the number of iterations $t/s$ timing the number of samples $N_{s}$, which derives $\Theta(N_{s}t/s)$ [33]. In the numerical experiment setting of Section V.5.1, $s=\Delta t=O(\Delta x^{2})=O(1/n^{2})$, the query complexity becomes $\Theta(N_{s}tn^{2})$.
@@ B.2.4 Classical query complexity for NAGD
The classical query complexity for NAGD is same as SGD, which is $\Theta(N_{s}t/s)$. In the numerical experiment setting of Section V.5.2, it becomes $\Theta(N_{s}tn^{2})$.
@@ B.3 Comparison of query complexity for different algorithms
From the above discussion, when $t$ is significant, loosely speaking, QLD and QAA have similar quantum query complexity $\tilde{O}(t)$ which is much better than $\tilde{O}(t^{4})$ of QHD. Furthermore, when the total evolution time is kept consistent across all quantum and classical algorithms, the quantum query complexity of QLD becomes comparable to those of SGD and NAGD, which are also $\tilde{O}(t)$.
This also indicates the superiority of QLD in scenario where the evolution time is significant.

### v2

@@ 2. $f(T)\geq T$ for $\forall T>0$,
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}+\eta\cdot\left(\left\langle{H}\right\rangle_{t}-2kT(t)\right)\leq 0.$ (83)
Defining a function $h(t)=(\left\langle{H}\right\rangle_{t}-2kf(T(t)))e^{\eta t}$, its time derivative can be calculated as
$\frac{\mathrm{d}h(t)}{\mathrm{d}t}=$ $\left(\frac{\mathrm{d}}{\mathrm{d}t}\left\langle{H}\right\rangle_{t}-2k\frac{\mathrm{d}f}{\mathrm{d}T}\frac{\mathrm{d}T}{\mathrm{d}t}\right)e^{\eta t}+(\left\langle{H}\right\rangle_{t}-2kf(T(t)))\eta e^{\eta t}.$ (84)
@@
$\left(\frac{\mathrm{d}}{\mathrm{d}t}\left\langle{H}\right\rangle_{t}-2k\frac{\mathrm{d}f}{\mathrm{d}T}\frac{\mathrm{d}T}{\mathrm{d}t}\right)+\eta(\left\langle{H}\right\rangle_{t}-2kf(T(t)))$ (86)
$\leq$ $\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}+\eta\cdot\left(\left\langle{H}\right\rangle_{t}-2kT(t)\right)\leq 0,$ (87)
@@
Then $h(t)\leq h(0)$ implies $\left\langle{H}\right\rangle_{t}-2kf(T(t))\leq O(e^{-\eta t})$. Using Eq. (23), we get $\left\langle{V}\right\rangle_{t}\leq\left\langle{H}\right\rangle_{t}\leq 2kf(T(t))+O(e^{-\eta t})$.
@@
5.1.4 Three time-dependent parameters $\eta(t),\hbar(t),T(t)$
@@
whose proof, similar to the proof of Corollaries 1, 2 and 3, is omitted here. The proof for quasar-convex case is essentially the same, where the convergence still holds
[…]
5.2 Objective functions
@@
Ref. [53] also explains how to categorize different objective functions. The category “Ridges or Valleys” is featured by small deep sections that divide the domain. “Basin” is featured by flatness near the global minimum of function values. “Flat” is featured by its flatness at elevated function values. “Studded” is identified by a foundational shape that exhibits high-frequency perturbations. The category “Simple” are functions that can be solved efficiently by standard gradient descent algorithms. The functions in each category are listed in Table 1, whose precise formulations can be found in Appendix A.2.
@@
5.3 Time-dependent Quantum Langevin Dynamics
In the previous experiments, we analyze roles of $T,\eta,\hbar$ in QLD, which also lead to the question why QLD would have a better performance if these parameters are time-dependent, as different ranges of parameters have different effect in different phases of the optimization process. At the beginning of the evolution, we set a high temperature to prompt the system to escape from local minimum. Then, we reduce the temperature to cool the system, allowing it to reach a more optimal, lower energy state. The discussion concerning $\hbar$ is also similar. At the beginning of the evolution, larger $\hbar$ can accelerate the rate of quantum tunneling, helping system escape from local minimum. After seeking the global minimum, smaller $\hbar$, weakening the quantum tunneling effect, can prevent system from tunneling back to local minimum. If $\hbar$ and $T$ decrease at the same time, $\hbar$ must decrease faster than $T$ as the effect of tunneling back to local minimum is very significant at a mild temperature, termed as “quantum reverse tunneling”. This could be due to the large value of $\hbar$, which violates the approximation $\hbar\Omega\ll kT$ as stated in Corollaries 1, 2 and 3. This violation could trigger the “quantum reverse tunneling” phenomenon. Therefore, $\hbar(t)$ needs to decrease faster than $T(t)$ as Eq. (98) and Eq. (99). This might ensure that the system can concentrate on the global minimum after escaping from local minimum.
@@ 5.3 Time-dependent Quantum Langevin Dynamics
Same finite difference method and implicit Euler method as Section 4.2 are used to numerically solve the evolution of system. The only difference is that time-dependent $\hbar(t)$ and $T(t)$ make $\hat{\mathcal{L}}$ of Eq. (105) be a time-dependent matrix ^$\hat{\mathcal{L}}(t)$. Dirichlet boundary condition and approximately uniform distribution are still kept here.
@@ 5.4 Quantum algorithms
5.4.1 Quantum Hamiltonian Descent (QHD)
@@
Then, we discretize the space into grids similar to Section 4.2 and use symplectic leapfrog scheme [27] to evolve this Hamiltonian system.
@@
5.4.2 Quantum Adiabatic Algorithm (QAA)
@@ 5.5 Classical algorithms
5.5.1 Stochastic Gradient Descent (SGD)
@@
5.5.2 Nesterov’s accelerated gradient descent (NAGD)
@@ 6 Discussion and Conclusions
However, there are several issues that necessitate further investigation. Theoretically, we believe that both the advantages and shortcomings of QLD can be interpreted more thoroughly.
In particular, due to the complex formula of the quantum Langevin dynamics, we only provide the convergence of QLD under the approximations of $\hbar\Omega\ll kT$ and $T\to 0$. Since the limit $\hbar\to 0$ clearly gives a transition to the classical Langevin equation, where all commutators vanish in this limit, in a sense, we might just prove the convergence of QLD in the classical limit, and use its quantum effect in the process of convergence. Whether the convergence still exists without these approximations remains further discussion with more meticulous mathematical and physical analysis. Numerically, since the simulation in the one-dimensional landscape alone has consumed a considerable amount of our computational resources, numerical experiments in high dimension requires more efficient numerical methods and more powerful classical computers. This also demonstrates the necessity of analog (or digital) quantum computers.
Acknowledgements
We thank Zhiyan Ding, Jiaqi Leng, Lin Lin, Hao Wu, Yukai Wu, Shuohan Zhang, and Zhennan Zhou for helpful discussions. Z. C. and Y. L. acknowledge research funding from Tsien Excellence in Engineering Program of Tsinghua University and Tsinghua University Initiative Scientific Research Program. T. L. was supported by the National Natural Science Foundation of China (Grant Numbers 62372006 and 92365117), and a startup fund from Peking University.
@@ A.1 Implementation details of numerical experiments for QLD
Because the objective function reaches positive infinity at the boundary of the domain in most cases, it is reasonable to use the Dirichlet boundary condition in our numerical experiments. However, other efforts are still needed to avoid the instability of the numerical calculation. In order to maintain the stability at the boundary, we set initial probability distribution to be an approximately uniform distribution with form
@@ A.2 The precise formulations of objective functions
We choose seven objective functions with different features in Table 1 to test the performance of different algorithms, as discussed in Section 5.2. The precise formulations and global minimum $x^{*}$ of these objective functions defined in $x\in[-1,1]$ are listed below,
@@ A.3 Analysis of numerical results
The results of Michalewicz function are shown in Section 5.6. Due to space constraints, the numerical results of rest six objective functions are shown here. The success probabilities of six objective functions for different algorithms at different (effective) evolution times are shown in Fig. 19. The probability distributions of six objective functions for different algorithms at different (effective) evolution times are shown in Figs. 20, 21, 22, 23, 24 and 25.
Based on the observations of Fig. 19, QLD exhibits much higher success probability across most objective functions than other algorithms. Specifically, the phenomena of strong quantum tunneling and thermal effects equip QLD with superior capabilities to escape from local minima encircled by numerous shallow wells at the beginning of evolution. This advantage enables QLD to achieve nearly 100% success probability in the Michalewicz, Alpine 1, Bohachevsky 2, and Deflected Corrugated Spring (DCS) functions which own several not-so-deep wells. However, the persistence of quantum noise and thermal effects towards the end of the evolution process hinder QLD’s ability to focus on the global minimum. Consequently, QLD performs poorly on the “Basin” function (Csendes), characterized by a flat region near the global minimum. The observations of dynamical processes in these experiments further confirms the three-phase explanation of QLD, which is shown at the end of Section 5.6.
@@ B.1 Quantum query complexity for QLD
We can divide Hamiltonian into two parts $H=E_{k}+V$. Then by finite difference method mentioned in Section 4.2, it is very easy to present $E_{k}$ and $V$ in the matrix forms. Consider discretizing the domain into $n$ grids along each axis, $E_{k}=-\frac{\hbar^{2}\mathbf{L}}{2m\Delta x^{2}}$, where $\mathbf{L}$ is the Laplacian matrix associated to the central difference method and $\Delta x$ is the width of discretizing grid [59], and $V=\text{diag}(V(x_{1}),\ldots,V(x_{n}))$ is a diagonal matrix by discretizing $V(x)$ in the grid, which is given by the oracle $O_{V}$ [53].
@@
Definition 3 (Block-encoding, Definition 43 of [34]).
@@
Definition 4 (State preparation pair, Definition 51 of [34]).
@@
Proposition 3 (Linear combination of block-encoded matrices, Proposition 5 of [74]).
@@
$\tilde{W}_{A}=$ $\big{(}P_{L}^{\dagger}\otimes I_{a+q}\big{)}\left(\sum_{j=0}^{m-1}\ket{j}\bra{j}\otimes(I_{a-a_{j}}\otimes U_{j})+\sum_{j=m}^{2^{b}-1}\ket{j}\bra{j}\otimes I_{a+q}\right)\big{(}P_{R}\otimes I_{a+q}\big{)}$
@@
Proposition 4 (Complexity for Lindblad simulation, Theorem 11 of [57]).
@@ B.2.3 Classical query complexity for SGD
For SGD, we assume that the classical evaluation oracle $O_{V}^{cl}$ allows us to query the function value of $V$ in a local way, so the total number of queries to $O_{V}^{cl}$ is at least the number of iterations $t/s$ timing the number of samples $N_{s}$, which derives $\Theta(N_{s}t/s)$ [59]. In the numerical experiment setting of Section 5.5.1, $s=\Delta t=O(\Delta x^{2})=O(1/n^{2})$, the query complexity becomes $\Theta(N_{s}tn^{2})$.
@@ B.2.4 Classical query complexity for NAGD
The classical query complexity for NAGD is same as SGD, which is $\Theta(N_{s}t/s)$. In the numerical experiment setting of Section 5.5.2, it becomes $\Theta(N_{s}tn^{2})$.
@@ B.3 Comparison of query complexity for different algorithms
From the above discussion, when $t$ is significant, loosely speaking, QLD and QAA have similar quantum query complexity $\tilde{O}(t)$ which is much better than $\tilde{O}(t^{4})$ of QHD. Furthermore, when the total evolution time is kept consistent across all quantum and classical algorithms, the quantum query complexity of QLD becomes comparable to those of SGD and NAGD, which are also $\tilde{O}(t)$. This also indicates the superiority of QLD in scenario where the evolution time is significant.